A committee of 5 is to be selected from a group of 6 men and 9 women. If the selection is made randomly, what is the probability that the committee consists of 3 men and 2 women?


1
Expert's answer
2021-10-25T17:41:45-0400

"P=\frac{C_6^3*C_9^2}{C_{15}^5}=\frac{20*36}{3003}=0.2398."